# A heralded single-photon source based on superpositions of squeezed states

### v1

@@ Abstract
We propose a heralded single-photon source based on injecting a superposition of oppositely squeezed states onto a beam splitter. Our superposition of squeezed states is composed of only even photon number states (the number of photons is equal to $2,6,10,...$) meaning the probability for an emitted single photon given a heralded singe photon events is higher than what one can achieve from the usual two-mode squeezed state. This enables one to realize an enhanced heralded single-photon source. We discuss how to create this superposition of squeezed states utilizing single-mode squeezed states and the cross-Kerr nonlinearity. Our proposed method significantly improves the probability of emitting the heralded single photon compared to spontaneous parametric down conversion.
@@ I Introduction
Single photon sources are essential tools in many quantum information-based technologies including quantum key distribution protocols [1] and linear optical quantum computation [2, 3]. As such there has been significant world-wide effort to propose and demonstrate on-demand single photon emitters. Quantum dots [4, 5, 6, 7, 8], nitrogen-vacancy centers in diamond [9, 10, 11], and negative silicon vacancy centers in diamond [12, 13, 14] have been proposed as solid state devices for realizing such emitter. Further cavity QED systems have been considered as a triggered single-photon source [15, 16, 17, 18] but the current work horse is still spontaneous parametric down conversion (SPDC) where we trigger off the idler photon [19, 20, 21]. However, the photon pair emission rate is low [$4\times 10^{-6}$ per pump photon using a periodically poled lithium niobate (PPLN) waveguide [22]].
@@
The essence of the heralded single-photon source is the generation of pairs of photons and the reason why we use SPDC. A more efficient method for generating entangled photons is the injection of squeezed light into a 50-50 beam splitter [23]. The resulting two-mode squeezed state source can in principle generate pairs of photons with a much higher success probability. There is however still room for improvement.
@@
In this paper, we introduce a new and more efficient approach based on superposition of squeezed state. We investigate how to generate a pair of photons by injecting an odd superposition of squeezed states onto a 50-50 beam splitter [24]. In particular we show that we can create this superposition using a single-mode squeezed state source and a cross-Kerr nonlinear medium although its heralded success probability is less than $1/2$ [25]. Finally, we determine the second-order intensity correlation functions $g^{(2)}(0)$ for the single-photon emissions of the superposition of oppositely squeezed states and the two-mode squeezed state. We use the measure $g^{(2)}(0)$ to evaluate whether or not the physical properties of a given photon source are close to those of an on-demand single-photon source [26].
@@
This article is organized as follows. In Sec. II, we begin by discussing the generation of a pair of photons using our superposition of oppositely squeezed states determining in Secs. II.1 and II.2 the probability and efficiency of emission of the heralded single photon source. Then Sec. II.3, we consider errors induced by imperfect operations of devices. Next in Sec. III, we compare our proposed heralded source with that from the pure two-mode squeezed state. Now in Sec. IV, we determine the second-order intensity correlation functions for our source using imperfect single-photon detectors. Finally in Sec. V we provide a brief concluding discussion.
@@ II Photon pair generation from a superposition of oppositely squeezed states
In our scheme, we assume that we have a stable source of squeezed states $|r\rangle$, as the generation of squeezed states is a mature technology. A typical squeezing level of squeezed light used for experiments is given by $6.3$ dB which corresponds to the squeezing parameter $|r|=0.725$ [27]. In the current paper, we typically use this parameter for showing examples but note that the detection of a $15$ dB squeezed state of light was demonstrated experimentally [28].
Figure 1: Schematic illustration of a heralded single photon source using squeezed state and a cross-Kerr medium.
II.1 Heralded single-photon probability and efficiency
@@
We can then define our generation efficiencies $E(r;\pm)$ and $E(r)$ as
$E(r;\pm)=P(1,1;r;\pm)/\sum_{n=0}^{\infty}P(1,n;r;\pm).$ (10)
$E(r)=P(1,1;r)/\sum_{n=0}^{\infty}P(1,n;r).$ (11)
If we observe |±$|\pm\alpha\rangle_{2}$ with the homodyne measurement, inject $|r;\pm\rangle_{a}$ into the beam splitter, and detect the heralding single photon in one port, we then obtain a single photon from the other port with probability $E(r;\pm)$. Similarly if we inject $|r\rangle_{a}$ into the beam splitter and detect the heralding single photon, we obtain a single photon with probability $E(r)$.
@@ II.2 Ideal single photon detection
In Fig. 2, we plot the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ that we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the input states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. Looking at Fig. 2, we note that $P(1,1;r;-)=0.453$, $P(1,5;r;-)=7.85\times 10^{-3}$, and $P(1,n;r;-)=0$ for $n=0,2,3,4$. Moreover, we obtain $E(r;-)=0.983$. If we generate the superposition of the oppositely squeezed states $|r;-\rangle_{a}$, it should allow us to realize the heralded single-photon source with a high probability and so be a good alternative to the SPDC. By contrast, if we prepare the squeezed state $|r\rangle_{a}$, we obtain $P(1,1;r)=7.54\times 10^{-2}$ with $E(r)=0.859$. Looking at the plot of $|r;-\rangle$, the following question comes to mind. There must be an optimal value of $r$ to maximize the probability. Answers to this question are shown in Fig. 3.
Figure 2: Plot of the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the injected states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. The dotted blue, solid red, and dashed purple line graphs represent $P(1,n;r)$, $P(1,n;r;-)$, and $P(1,n;r;+)$, respectively. The line graph of $P(1,n;r;+)$ overlaps the horizontal axis of $n$ mostly.
Figure 3: (a) Plot of the probability ${\cal N}_{-}(r)P(1,1;r;-)/4$ we obtain a heralded single photon by injecting $|r;-\rangle_{a}$ for $0\leq r\leq 2$. When $r=1.146$, the graph reaches its maximum value 0.096 230.096230.096{\,}230.096 23. Thus, the optimal value of $r$ to maximize the probability is $r=1.146$. (b) Plot of the efficiency and probabilities of the heralded single photon as functions of $r$. The thick solid red and thin solid blue curves represent $E(r;-)$ and $E(r)$, respectively. The thick dashed red and thin dashed blue curves represent $P(1,1;r;-)$ and $P(1,1;r)$, respectively.
@@
Next in Fig. 3(a), we plot the actual probability that we obtain the $|1\rangle$ photon Fock state conditioned that we observe a heralding single photon for the initial state $|r;-\rangle_{a}$. Examining Fig. 3(a), we note that ${\cal N}_{-}(r)P(1,1;r;-)/4$ reaches its maximum value of 0.096 230.096230.096{\,}230.096 23 when $r=1.146$. At that point, the efficiency is $E(1.146;-)=0.9488$. Figure 3(b) shows plots of the efficiencies and probabilities of the heralded single photon as a function of $r$. We observe that $E(r;-)$ is larger than or equal to $E(r)$ for any $r$. Thus, as the heralded single-photon source, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the typical squeezed state. It is useful to mention that the squeezed state $|r\rangle$ is a superposition of $\{|2n\rangle:n=0,1,2,...\}$ while contrastingly, the superpositions of the oppositely squeezed states $|r;+\rangle$ and $|r;-\rangle$ are superposition of $\{|2m\rangle:m=0,2,4,...\}$ and $\{|2m\rangle:m=1,3,5,...\}$, respectively. The Fock states present in these superpositions of squeezed states are further separated in photon number space than the usual squeezed state. If we can prepare $|r;-\rangle\propto c_{2}|2\rangle+c_{6}|6\rangle+...$ and inject that state into the 50-50 beam splitter, a transformation $|2\rangle_{a}|0\rangle_{b}\to(1/2)(|2\rangle_{a}|0\rangle_{b}+|0\rangle_{a}|2\rangle_{b})-(1/\sqrt{2})|1\rangle_{a}|1\rangle_{b}$ occurs and we obtain a state $|1\rangle_{a}|1\rangle_{b}$ with a maximal probability of $1/2$. Hence, we can implement the heralded single-photon source with higher efficiency. Figures 4(a) and (b) show 3D plots of the probabilities of generating $|r;-\rangle$ and a heralded single photon as functions of $\tilde{\tau}=2\kappa\tau$ and $r$, respectively, with fixing $\alpha$ at a specific value. In Fig. 4(a), we plot ${\cal P}_{0}$, the probability of generating $|r;-\rangle$, where
@@
Looking at both of the plots, we observe sharp peaks at $\tilde{\tau}=\pi$.
@@ II.3 Imperfections and errors
In any realistic implementation, there will be imperfections and so let us explore effects associated with the cross-phase modulation and also single-photon heralding detection. We have previously detailed the generation of $|r;-\rangle$ from the squeezed state $|r\rangle$ using the nonlinear cross-Kerr medium. For this generation, we had set duration $\tau$ for the nonlinear interaction as $2\kappa\tau=\pi$. Here, we consider a case where an inaccuracy occurs and such timing is not perfect and a phase error could be induced as $2\kappa\tau=\pi+\Delta\theta$. In this case
@@
and so evaluate the effects caused by $\Delta\theta$, we define the following ratio that implies the generation rate of $|r;-\rangle$:
@@
Here $\Delta\theta$ is not a constant value but a probabilistic random variable that obeys Gaussian distribution whose mean value and standard deviation are given by zero and $\sigma$, respectively. Taking the average of $R(r,\alpha,\Delta\theta)$ with varying $\Delta\theta$ for many samples, we obtain $R(r,\alpha,\sigma)$. In Fig. 5(a), we plot $R(r,\alpha,\sigma)$ for $\alpha=9$, $10$, and $11$ with $r=0.725$. In the range of $0\leq\sigma\leq 0.001$, the function $\exp(-\lambda\sigma^{2})$ approximate to the curves of $R(r,\alpha,\sigma)$ well. Looking at those plots, we note that the curve of $R(r,\alpha,\sigma)$ gets closer to a horizontal as $\alpha$ becomes smaller. Thus, we become aware that the error of the inaccuracy of the duration is serious for large $\alpha$. Because the homodyne measurement requires a strong coherent state, we need to increase $|\alpha|$ for precise measurements. However, large $|\alpha|$ makes the single-photon emission vulnerable to the inaccuracy of duration $\tau$ and a trade-off emerges. Figure 5(b) shows a 3D plot of $R(r,\alpha,\sigma)$ with $\alpha=10$ as a function of $r$ and $\sigma$ where we note that, in the range where $r$ is larger than $1/2$, $R(r,10,\sigma)$ decreases as $\sigma$ gets larger. Fluctuations of the phase error are problematic when we try to obtain strong squeezing. If we want to generate a strongly squeezed state, we must ensure the accuracy of the cross-Kerr phase modulation giving us another trade-off.
@@
$Q(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}\sum_{m_{b}=0}^{\infty}(1-\eta)^{n-1}P(n_{a},m_{b};r;-).$ (18)
@@
$Q_{1}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}(1-\eta)^{n-1}P(n_{a},1;r;-).$ (19)
Then, we can compute the efficiency as
$E(r,\eta;-)=\frac{Q_{1}(r,\eta;-)}{Q(r,\eta;-)}.$ (20)
Figure 6: (a) Plot of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as a function of $r$ for $\eta=0.9$, where they are represented by the thick solid red, thin solid blue, and thick dashed red curves, respectively. (b) Plot of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as a function of $r$ with $\eta=0.9$, where they are represented by the thick red and thin blue curves, respectively.
Figure 7: (a) 3D plots of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where they are represented by the curved red, blue, and green transparent surfaces, respectively. (b) 3D plots of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$, where they are represented by the curved red and blue transparent surfaces, respectively. In the regime $r\leq 2$ and $0.7\leq\eta\leq 1$ we have $Q_{1}(r,\eta;-)>\tilde{Q}_{1}(r,\eta)>{\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ and $E(r,\eta;-)<\tilde{E}(r,\eta)$.
In Fig. 6(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ for $\eta=0.9$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. Although the single-photon detector is imperfect ($\eta=0.9$), $Q_{1}(r,\eta;-)$ is nearly equal to $0.5$ for the small squeezing parameter $r$. This is one of the advantages that our method has.
Now in Fig. 6(b), we plot $E(r,\eta;-)$ as a function of $r$ for $\eta=0.9$. This plot shows that $E(r,\eta;-)$ is nearly equal to $0.95$ at $r=0$ and decreases gradually as $r$ becomes larger. In Fig. 7(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. In the limit of $r\to 0$, $Q_{1}(r,\eta;-)$ is finite and non-zero but ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ approaches zero. In Fig. 7(b), we plot $E(r,\eta;-)$ as a function of $r$ and $\eta$. As $r$ increases and $\eta$ decreases, $E(r,\eta;-)$ becomes smaller gradually.

### v2

# Heralded single-photon source based on superpositions of squeezed states

@@ Abstract
We propose a heralded single-photon source based on injecting a superposition of oppositely squeezed states onto a beam splitter. Our superposition of squeezed states is composed of only even photon number states (the number of photons is equal to $2,6,10,...$) meaning the probability for an emitted single photon given as a heralded single-photon event is higher than what one can achieve from the usual two-mode squeezed state. This enables one to realize an enhanced heralded single-photon source. We discuss how to create this superposition of squeezed states utilizing a single-mode squeezed state and the cross-Kerr nonlinearity. Our proposed method significantly improves the probability of emitting the heralded single photon compared to spontaneous parametric down-conversion.
@@ I Introduction
Single-photon sources are essential tools in many quantum information-based technologies including quantum key distribution protocols [1] and linear optical quantum computation [2, 3]. As such there has been significant world-wide effort to propose and demonstrate on-demand single-photon emitters. Quantum dots [4, 5, 6, 7, 8], nitrogen-vacancy centers in diamond [9, 10, 11], and negative silicon vacancy centers in diamond [12, 13, 14] have been proposed as solid state devices for realizing such emitters. Further cavity QED systems have been considered as a triggered single-photon source [15, 16, 17, 18], but the current work horse is still spontaneous parametric down-conversion (SPDC) where we trigger off the idler photon [19, 20, 21, 22, 23, 24]. However, the photon pair emission rate is low [$4\times 10^{-6}$ per pump photon using a periodically poled lithium niobate (PPLN) waveguide [25]].
Implementation of heralded single-photon sources with the SPDC is state-of-the-art technology and is being studied with great effort currently [26]. A review of the heralded single-photon sources was given in [27, 28]. Improvement of the heralded single-photon source with photon-number-resolving detectors was examined in Refs. [29, 30]. Suppression of multiphoton events in a cavity-enhanced SPDC via the photon-blockade effect for the heralded single-photon source has been proposed [31], and a technique of heralding based on the detection of zero photons was investigated [32].
@@
The essence of the heralded single-photon source is the generation of pairs of photons and is the reason we use SPDC. A more efficient method for generating entangled photons is the injection of squeezed light into a 50-50 beam splitter [33]. The resulting two-mode squeezed state source can in principle generate pairs of photons with a much higher success probability. Using an imperfect single-photon detector whose efficiency is equal to $0.9$, the pure two-mode squeezed state with a squeezing parameter $0.5$ gives the photon pair emission rate $0.151$. There is, however, still room for improvement.
@@
In this paper, we introduce a more efficient approach based on superpositions of squeezed states. We investigate how to generate a pair of photons by injecting an odd superposition of squeezed states onto a 50-50 beam splitter [34]. In particular we show that we can create this superposition using a single-mode squeezed state source and a cross-Kerr nonlinear medium, although its heralded success probability is less than $1/2$ [35, 36]. Finally, we determine the second-order intensity correlation functions $g^{(2)}(0)$ for the single-photon emissions of the superposition of oppositely squeezed states and the two-mode squeezed state. We use the measure $g^{(2)}(0)$ to evaluate whether or not the physical properties of a given photon source are close to those of an on-demand single-photon source [37].
@@
This article is organized as follows. In Sec. II, we begin by discussing the generation of a pair of photons using our superposition of oppositely squeezed states determining in Secs. II.1 and II.2 the probability and conditional probability of emission of the heralded single-photon source. Then in Sec. II.3, we consider errors induced by imperfect operations of devices. Next in Sec. III, we compare our proposed heralded source with that from the pure two-mode squeezed state. In Sec. IV, we determine the second-order intensity correlation functions for our source using imperfect single-photon detectors. Finally in Sec. V we provide a brief concluding discussion.
@@ II Photon pair generation from a superposition of oppositely squeezed states
In our scheme, we assume that we have a stable source of squeezed states $|r\rangle$, as the generation of squeezed states is a mature technology. A typical squeezing level of squeezed light used for experiments is given by $6.3$ dB, which corresponds to the squeezing parameter $|r|=0.725$ [38]. In the current paper, we typically use this parameter for showing examples but note that the detection of a $15$ dB squeezed state of light was demonstrated experimentally [39].
In the proposed method, it is critical to adjust the duration of the nonlinear interaction time $\tau$. To realize a precise adjustment of $\tau$, we can fabricate a cavity of the cross-Kerr nonlinear medium, which is caught between two half-silvered mirrors. We can vary $\tau$ by changing the transmittance of mirrors. To obtain the phase factor $e^{-i2n\kappa\tau}$with $2\kappa\tau=\pi$ for Eq. (3) precisely, $\kappa$ must be large enough.
Figure 1: Schematic illustration of a heralded single-photon source using a squeezed state and a cross-Kerr medium.
II.1 Heralded single-photon probability and conditional probability
@@
We can then define our conditional probabilities $P_{\mbox{\scriptsize c}}(r;\pm)$ and $P_{\mbox{\scriptsize c}}(r)$ as
$P_{\mbox{\scriptsize c}}(r;\pm)=P(1,1;r;\pm)/\sum_{n=0}^{\infty}P(1,n;r;\pm).$ (10)
$P_{\mbox{\scriptsize c}}(r)=P(1,1;r)/\sum_{n=0}^{\infty}P(1,n;r).$ (11)
If we observe |±$|\pm\alpha\rangle_{2}$ with the homodyne measurement, inject $|r;\pm\rangle_{a}$ into the beam splitter, and detect the heralding single photon in one port, we then obtain a single photon from the other port with probabilities $P_{\mbox{\scriptsize c}}(r;\pm)$. Thus, we can regard $P_{\mbox{\scriptsize c}}(r;\pm)$ as the conditional probabilities of single-photon emission after the postselection with the measurement outcomes on the other port. Similarly if we inject $|r\rangle_{a}$ into the beam splitter and detect the heralding single photon, we obtain a single photon with conditional probability $P_{\mbox{\scriptsize c}}(r)$.
@@ II.2 Ideal single photon detection
In Fig. 2, we plot the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ that we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the input states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. Looking at Fig. 2, we note that $P(1,1;r;-)=0.453$, $P(1,5;r;-)=7.85\times 10^{-3}$, and $P(1,n;r;-)=0$ for $n=0,2,3,4$. Moreover, we obtain $P_{\mbox{\scriptsize c}}(r;-)=0.983$. If we generate the superposition of the oppositely squeezed states $|r;-\rangle_{a}$, it should allow us to realize the heralded single-photon source with a high probability and so be a good alternative to the SPDC. By contrast, if we prepare the squeezed state $|r\rangle_{a}$, we obtain $P(1,1;r)=7.54\times 10^{-2}$ with $P_{\mbox{\scriptsize c}}(r)=0.859$. Looking at the plot of $|r;-\rangle$, the following question comes to mind. There must be an optimal value of $r$ to maximize the probability. Answers to this question are shown in Fig. 3.
Figure 2: Plot of the probability $P(1,n;r)$ and $P(1,n;r;\pm)$. We generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the injected states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. The dotted blue, solid red, and dashed purple line graphs represent $P(1,n;r)$, $P(1,n;r;-)$, and $P(1,n;r;+)$, respectively. The line graph of $P(1,n;r;+)$ overlaps the horizontal axis of $n$ mostly.
Figure 3: (a) Plot of the probability ${\cal N}_{-}(r)P(1,1;r;-)/4$. We obtain a heralded single photon by injecting $|r;-\rangle_{a}$ for $0\leq r\leq 2$. When $r=1.146$, the graph reaches its maximum value 0.096 230.096230.096{\,}230.096 23. Thus, the optimal value of $r$ to maximize the probability is $r=1.146$. (b) Plot of the probabilities of the heralded single photon as functions of $r$. The thick solid red and thin solid blue curves represent $P_{\mbox{\scriptsize c}}(r;-)$ and $P_{\mbox{\scriptsize c}}(r)$, respectively. The thick dashed red and thin dashed blue curves represent $P(1,1;r;-)$ and $P(1,1;r)$, respectively.
@@
Next in Fig. 3(a), we plot the actual probability that we obtain the $|1\rangle$ photon Fock state conditioned that we observe a heralding single photon for the initial state $|r;-\rangle_{a}$. Examining Fig. 3(a), we note that ${\cal N}_{-}(r)P(1,1;r;-)/4$ reaches its maximum value of 0.096 230.096230.096{\,}230.096 23 when $r=1.146$. At that point, the conditional probability is $P_{\mbox{\scriptsize c}}(1.146;-)=0.9488$. Figure 3(b) shows plots of the probabilities of the heralded single photon as a function of $r$. We observe that $P_{\mbox{\scriptsize c}}(r;-)$ is larger than or equal to $P_{\mbox{\scriptsize c}}(r)$ for any $r$. Thus, as the heralded single-photon source, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the typical squeezed state. It is useful to mention that the squeezed state $|r\rangle$ is a superposition of $\{|2n\rangle:n=0,1,2,...\}$ while contrastingly, the superpositions of the oppositely squeezed states $|r;+\rangle$ and $|r;-\rangle$ are superpositions of $\{|2m\rangle:m=0,2,4,...\}$ and $\{|2m\rangle:m=1,3,5,...\}$, respectively. The Fock states present in these superpositions of squeezed states are further separated in photon number space than the usual squeezed state. If we can prepare $|r;-\rangle\propto c_{2}|2\rangle+c_{6}|6\rangle+...$ and inject that state into the 50-50 beam splitter, a transformation $|2\rangle_{a}|0\rangle_{b}\to(1/2)(|2\rangle_{a}|0\rangle_{b}+|0\rangle_{a}|2\rangle_{b})-(1/\sqrt{2})|1\rangle_{a}|1\rangle_{b}$ occurs, and we obtain a state $|1\rangle_{a}|1\rangle_{b}$ with a maximal probability of $1/2$. Hence, we can implement the heralded single-photon source with higher conditional probability. Figures 4(a) and 4(b) show 3D plots of the probabilities of generating $|r;-\rangle$ and a heralded single photon as functions of $\tilde{\tau}=2\kappa\tau$ and $r$, respectively, with fixing $\alpha$ at a specific value. In Fig. 4(a), we plot ${\cal P}_{0}$, the probability of generating $|r;-\rangle$, where
@@
Looking at both plots, we observe sharp peaks at $\tilde{\tau}=\pi$.
@@ II.3 Imperfections and errors
In any realistic implementation, there will be imperfections, and so let us explore effects associated with the cross-phase modulation and also single-photon heralding detection. We have previously detailed the generation of $|r;-\rangle$ from the squeezed state $|r\rangle$ using the nonlinear cross-Kerr medium. For this generation, we had set duration $\tau$ for the nonlinear interaction as $2\kappa\tau=\pi$. Here, we consider a case where an inaccuracy occurs; such timing is not perfect and a phase error could be induced as $2\kappa\tau=\pi+\Delta\theta$. In this case
@@
and so to evaluate the effects caused by $\Delta\theta$, we define the following ratio that implies the decrease in generation probability of $|r;-\rangle$ due to $\Delta\theta$:
@@
Here $\Delta\theta$ is not a constant value but a probabilistic random variable that obeys Gaussian distribution whose mean value and standard deviation are given by zero and $\sigma$, respectively. Taking the average of $R(r,\alpha,\Delta\theta)$ with varying $\Delta\theta$ for many samples, we numerically obtain
$R(r,\alpha,\sigma)=\int_{-\infty}^{\infty}d(\Delta\theta)\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp(-\frac{\Delta\theta^{2}}{2\sigma^{2}})R(r,\alpha,\Delta\theta).$ (17)
In Fig. 5(a), we plot $R(r,\alpha,\sigma)$ for $\alpha=9$, $10$, and $11$ with $r=0.725$. In the range of $0\leq\sigma\leq 0.001$, the function $\exp(-\lambda\sigma^{2})$ approximates to the curves of $R(r,\alpha,\sigma)$ well. Looking at those plots, we note that the curve of $R(r,\alpha,\sigma)$ gets closer to a horizontal as $\alpha$ becomes smaller. Thus, we become aware that the error of the inaccuracy of the duration is serious for large $\alpha$. Because the homodyne measurement requires a strong coherent state, we need to increase $|\alpha|$ for precise measurements. However, large $|\alpha|$ makes the single-photon emission vulnerable to the inaccuracy of duration $\tau$, and a trade-off emerges. Figure 5(b) shows a 3D plot of $R(r,\alpha,\sigma)$ with $\alpha=10$ as a function of $r$ and $\sigma$ where we note that, in the range where $r$ is larger than $1/2$, $R(r,10,\sigma)$ decreases as $\sigma$ gets larger. Fluctuations of the phase error are problematic when we try to obtain strong squeezing. If we want to generate a strongly squeezed state, we must ensure the accuracy of the cross-Kerr phase modulation giving us another trade-off.
In Fig. 5(a), $R(r,\alpha,\sigma)$ departs from unity considerably for $\sigma=0.004$. This standard deviation corresponds to a fluctuation $\Delta\tau=\tau\sigma/\pi$. This evaluation suggests to us that we must fix $\tau$ with the accuracy $\Delta\tau/\tau\sim 10^{-3}$, and it is well within our reach experimentally.
@@
$P_{\mbox{\scriptsize click}}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}\sum_{m_{b}=0}^{\infty}(1-\eta)^{n_{a}-1}P(n_{a},m_{b};r;-).$ (19)
@@
$P_{\mbox{\scriptsize click},1}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}(1-\eta)^{n_{a}-1}P(n_{a},1;r;-).$ (20)
Then, we can compute the conditional probability as
$P_{\mbox{\scriptsize click,c}}(r,\eta;-)=\frac{P_{\mbox{\scriptsize click},1}(r,\eta;-)}{P_{\mbox{\scriptsize click}}(r,\eta;-)}.$ (21)
Figure 6: (a) Plot of $P_{\mbox{\scriptsize click},1}(r,\eta;-)$, $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ defined in Eq. (25), and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as a function of $r$ for $\eta=0.9$, where they are represented by the thick solid red, thin solid blue, and thick dashed red curves, respectively. (b) Plot of $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ defined in Eq. (26) as a function of $r$ with $\eta=0.9$, where they are represented by the thick red and thin blue curves, respectively.
Figure 7: (a) 3D plots of $P_{\mbox{\scriptsize click},1}(r,\eta;-)$, $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$, and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where they are represented by the curved red, blue, and green transparent surfaces, respectively. (b) 3D plots of $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ and $\eta$, where they are represented by the curved red and blue transparent surfaces, respectively. In the regime $r\leq 2$ and $0.7\leq\eta\leq 1$ we have $P_{\mbox{\scriptsize click},1}(r,\eta;-)>\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)>{\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ and $P_{\mbox{\scriptsize click,c}}(r,\eta;-)<\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$.
In Fig. 6(a), we plot $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ for $\eta=0.9$, where ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. Although the single-photon detector is imperfect ($\eta=0.9$), $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is nearly equal to $0.5$ for the small squeezing parameter $r$. This is one of the advantages that our method has.
In Fig. 6(b), we plot $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ as a function of $r$ for $\eta=0.9$. This plot shows that $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ is nearly equal to $0.95$ at $r=0$ and decreases gradually as $r$ becomes larger. In Fig. 7(a), we plot $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. In the limit of $r\to 0$, $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is finite and nonzero, but ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ approaches zero. In Fig. 7(b), we plot $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ as a function of $r$ and $\eta$. As $r$ increases and $\eta$ decreases, $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ becomes smaller gradually.